OpenStudy (anonymous):

given f(x)=2x-2/4 solve for f-1(3)

OpenStudy (anonymous):

So assume y = f(x) = 2x - 2/4

OpenStudy (anonymous):

So what will x be in terms of y?

OpenStudy (anonymous):

so i swich x and y around right?

OpenStudy (anonymous):

not exactly, as y=2x-(2/4)

OpenStudy (anonymous):

y+(2/4) = 2x

OpenStudy (anonymous):

So what will be x?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

Nope x = (y-(2/4))/2

OpenStudy (anonymous):

Now plug in y = 3

OpenStudy (anonymous):

You will get the value of x which is f^-1(3)

OpenStudy (anonymous):

so x =3-(2/4)/2

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

how do i solve that?

OpenStudy (anonymous):

well it is (3-(2/4))/2

OpenStudy (anonymous):

so i.e (3-0.5)/2

OpenStudy (anonymous):

1.25

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thats not a choice i have

OpenStudy (anonymous):

Sorry I made a mistake in the signs earlier

OpenStudy (anonymous):



OpenStudy (anonymous):

1.75?

OpenStudy (anonymous):

yup